OpenStudy (anonymous):

An electron (mass: 9.11 × 10-31 kg) and a proton (mass: 1.67 × 10-27 kg) are each accelerated from rest through a potential difference of 15 V that spans 15 cm. Which one takes less time to cover the 15 cm distance? The proton beats the electron. It's a tie—each reaches 15cm at the same time. There isn't enough information given to answer this question. The electron beats the proton

OpenStudy (anonymous):

If I remember correctly, the electron is faster because if you both accelerate them through a potential difference of 15V, both their energies are 15eV. We are concerned with kinetic energy, given by .5mv^2. So we have 15eV=.5mv^2, which equals 30eV=mv^2, and to find V^2 we have to divide by m, or mass. And the smaller the mass, the larger the velocity, therefore the electron should be faster. tl;dr Both have same energy, the one with less mass has a higher velocity, given by 1/2mv^2, electron is lighter than proton, so it has the higher velocity.

OpenStudy (anonymous):

\[Fe=me*ae \] \[Fp= mpro*ap\] since they have the same charge \[me*ae=mpro*ap \rightarrow ae=(mpro/me)*ap>ap\] so eletron reach first

OpenStudy (anonymous):

Thanks:)